OpenStudy (anonymous):

Solve dy/dx=(-5x+y)^2 -4 I got 1/6ln((u-3)/(u+3))=x using a substitution of u=(-5x+y). Is that right?

OpenStudy (anonymous):

the problem isnt stated properly... it looks wrong though. some clarification would help.

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

\[dy/dx=(-5x+y)^2-4\] Is that better?

OpenStudy (anonymous):

That u-substitution works for this differential equation. The substitution reduces it to the variables-separable equation \[du/(u^{2}-9) + dx = 0.\]

OpenStudy (anonymous):

But then you still have to integrate it, right?

OpenStudy (anonymous):

Did you get the same answer I got?

OpenStudy (anonymous):

Computing ...

OpenStudy (anonymous):

(1/6)(ln|u - 3| - ln|u + 3| = -x + ln(C)/6 as an intermediate solution. Once that is taken back to original variables, we're finished.

OpenStudy (anonymous):

My final solution is 5x - y + 3 = C e^(6x)(5x - y - 3) in implicit form.

OpenStudy (anonymous):

Its explicit form is \[y = 5 x + 3(1 + C e^{6x})/(-1 + C e^{6x}).\]

OpenStudy (anonymous):

For increasingly large values of x, all solution curves more closely approximate the linear function 5x + 3, so the line y = 5x + 3 is the asymtote of all solution curves.

OpenStudy (anonymous):

That's a good problem, Dave :^)

OpenStudy (anonymous):

I'll tell my professor;-) Thanks!

OpenStudy (anonymous):

Major ten-four, good buddy :^)